Vanadium atom has a radius of 121 pm and crystallizes with a BCC unit cell. Determine the number of unit cells present in 1.9 cm3 solid sample of Vanadium. Express answer in scientific notation


1
Expert's answer
2021-10-07T03:44:15-0400

Radius=121 pm "= 121 \times 10^{-8}"

BBC No of atoms in unit cell = 2

in BBC

"a=\frac{4}{\sqrt{3}} \times" (Relation between edge length and atom radius)

"a=\frac{4}{\sqrt{3}}\times 1.21 \times 10^{-8} \\

a=2.794 \times 10^{-8}\;cm"

Volume of one unit cell "= a^3 =(2.794 \times 10^{-8})^3 = 2.181 \times 10^{-23} \;cm^3"

No of unit cells present in "1.9 \;cm^3"

"=\frac{1.9}{Volume \; of \; one \; unit \;cell} \\

= \frac{1.9}{2.181 \times 10^{-23}} \\

= 8.711 \times 10^{22}"

No of unit all present in "1.9\; cm^3 = 8.711 \times 10^{22}" unit cells.